# The Cross-Lingual Cost: Retrieval Biases in RAG over Arabic-English Corpora

### v1

@@ 4.4 Mitigating Cross-lingual Failings
These results raise an important question: can multilingual retrievers be used reliably on mixed-language corpora without further tuning? To explore this question, we introduce the balanced retriever, which enforces an equal selection of documents from each language by retrieving 10 passages in Arabic and 10 in English. We evaluate this approach in the same experimental setup described earlier. The third and final bar in each subfigure of Figure 1 presents its results. Notice that while the language-oracle retriever is infeasible, it serves as an upper bound for what the balanced retrieval could achieve. In practice, we observe that the balanced retriever does not significantly degrade performance on same-language cases while providing substantial improvements in cross-lingual cases. Notably, the balanced retriever yields more consistent results across the different combinations of user and document languages, unlike the direct setting, which favors the same-language combinations at the expense of cross-language ones. Moreover, this strategy leads to a significant improvement in overall retrieval accuracy across benchmarks and embedders, with consistent gains of around 3-5% for BGE-M3 and approximately 20% for M-E5.
These findings suggest that even simple strategies can help mitigate retrieval biases, indicating that debiasing is feasible despite inherent biases in the embedding model.
@@ 5 Conclusions
Our analysis shows that performance degrades most in cross-lingual settings where the user and document languages differ, with drops of 30–50% compared to same-language configurations. Using an oracle retriever restricted to the correct language, we isolate the primary source of failure as the retriever’s difficulty in ranking documents across languages. That is, while the retriever performs well within a single language, it struggles when comparing passages across languages, often favoring those in the query’s language. We further observe that different embedders exhibit weaknesses in different cross-lingual settings. This highlights the potential to improve training by explicitly targeting cross-lingual robustness and narrowing the gap with same-language performance.
Lastly, we demonstrate that a simple mitigation of retrieving a balanced number of documents per language, can significantly improve cross-lingual performance and even improve overall results. This finding points to meaningful opportunities for mitigating multilingual retrieval biases, particularly in real-world applications. However, applying such a balancing approach in practical settings with non-uniform language distributions or more than two languages remains an open challenge and requires further investigation.
@@ Acknowledgments
We thank our colleagues at AI71, and in particular Abdelrahman Ibrahim, Amr Ali Abugreedah, Mohamad Salah, Saleem Hamo, Kirollos Sorour, Imran Moqbel, and Anas AlHelali, whose native proficiency in Arabic was instrumental in the annotation process used to validate the evaluation metrics in our pipeline.
@@ A.1.1 Answer Accuracy
[…]
@@ A.1.2 Retrieval Hit@20
Now that we trusted our LLM-based accuracy metric, we moved to validating whether our Hits@20 metric, which also uses LLM judgments, effectively captures success in the retrieval step. Towards this goal, we analyzed the downstream accuracy as a function of the Hit@20 score. This analysis focused on the Legal benchmark, where the no-RAG accuracy is relatively low (27%), making it easier to observe the impact of retrieval quality. Table 3 reports these results for the BGE-M3 and M-E5 embedders.

### v2

@@ 4.4 Mitigating Cross-lingual Failings
These results raise an important question: can multilingual retrievers be used reliably on mixed-language corpora without further tuning? To address this question, we consider two retrieval baselines. The first, denoted translation, translates each query into the other language using the Google Translate API and performs retrieval twice (once per language, since the document language is not known a priori). The two ranked lists are then merged, and the top 20 results are selected according to the retrieval score (i.e., the inner product of the query and document embeddings). The second method, balanced, enforces equal selection across languages by retrieving 10 passages in Arabic and 10 in English.
We evaluate these approaches under the same experimental setup described earlier. The last two bars in each subfigure of Figure 1 present the corresponding results. While the language-oracle retriever is not feasible in practice, it serves as an upper bound for what the translation and balanced approaches could achieve. In practice, both the translation and balanced retrievers show no statistically significant loss relative to the direct retriever in same-language cases, while providing substantial improvements in cross-lingual cases. Notably, those retrievers yields more consistent results across the different combinations of user and document languages, unlike the direct setting, which favors the same-language combinations at the expense of cross-language ones. Moreover, this strategy leads to a considerable improvement in overall retrieval accuracy across benchmarks and embedders, with consistent gains of around 4-6% for BGE-M3 and approximately 20% for M-E5.
No statistically significant difference is observed between the performance of the translation retriever and the balanced retriever. However, they differ in latency and cost: while translation requires an expensive and time-consuming call to a translation service, balanced incurs no additional cost beyond retrieving documents from the index. One might also assume that balanced requires prior knowledge of the proportion of ground-truth documents in each language. Yet, as shown in Appendix A.5, the performance of balanced remains stable even when this proportion varies, suggesting that balanced may offer a more practical alternative in such scenarios.
@@ 5 Conclusions
Our analysis shows that performance degrades most in cross-lingual settings where the user and document languages differ, with drops that can exceed 40% compared to same-language configurations. Using an oracle retriever restricted to the correct language, we isolate the primary source of failure as the retriever’s difficulty in ranking documents across languages. That is, while the retriever performs well within a single language, it struggles when comparing passages across languages, often favoring those in the query’s language. We further observe that different embedders exhibit weaknesses in different cross-lingual settings. This highlights the potential to improve training by explicitly targeting cross-lingual robustness and narrowing the gap with same-language performance.
Lastly, we show that simple mitigations, such as retrieving a balanced number of documents per language or translating the query, can substantially improve cross-lingual performance and even enhance overall results. This finding highlights meaningful opportunities for reducing multilingual retrieval biases, particularly in real-world applications. However, applying such approaches in practical settings with non-uniform language distributions or more than two languages remains an open challenge and warrants further investigation.
@@ Acknowledgments
We thank our colleagues at AI71, and in particular Abdelrahman Ibrahim, Amr Ali Abugreedah, Mohamad Salah, Saleem Hamo, Kirollos Sorour, Imran Moqbel, and Anas AlHelali, whose native proficiency in Arabic was instrumental in the annotation process used to validate the evaluation metrics in our pipeline, as well as Michal Caspi for co-leading this effort.
@@ A.1.1 Answer Accuracy
Independently, human annotators who are native speakers of the respective languages were asked to assess the similarity between the generated and reference answers. The annotators were asked to label each pair as matching or not matching and to mark them as debatable or non-debatable. Of the non-debatable items (80% in both languages) the agreement rate was 95% for English and 98% for Arabic. The overall agreement rates are 82% for English and 85% for Arabic, which means that almost all disagreements were for cases marked as debatable. Therefore, the annotations corroborate the validity of the automated accuracy metric.
@@ A.1.2 Retrieval Hit@20
Now that we trusted our LLM-based accuracy metric, we moved to validating whether our Hits@20 metric, which also uses LLM judgments, effectively captures success in the retrieval step. Toward this goal, we analyzed the downstream accuracy as a function of the Hit@20 score. This analysis focused on the Legal benchmark, where the no-RAG accuracy is relatively low (27%), making it easier to observe the impact of retrieval quality. Table 3 reports these results for the BGE-M3 and M-E5 embedders.
@@ A.2 Benchmark configuration and statistics
The benchmark was balanced after the DataMorgana filtering step to include 50% questions grounded in Arabic documents and 50% in English documents. Statistics for the final benchmark are presented in Table 4.
Benchmark
Query
language
Document
language
Count
Legal English English 318
English Arabic 337
Arabic English 324
Arabic Arabic 303
Travel English English 513
English Arabic 471
Arabic English 479
Arabic Arabic 460
Table 4: Benchmark breakdown by query and document language.
@@ A.3 Prompts
A.4 Additional Results with NDCG and MRR
Table 5 provides the counterpart to Table 2, augmented with MRR@20 and NDCG@20 results. Figures 2 and 3 present a variation of Figure 1 but according to NDCG@20 and MRR@20 respectively. The results for the balanced retriever are omitted because ranking-based metrics like NDCG@20 and MRR@20 require a single, consistent ordering of retrieved passages. Since the balanced retriever returns two separate rank lists (one in Arabic and one in English), these metrics cannot be meaningfully computed. As can be seen, consistent trends occur for all metrics: Cross-language performance is worse compared to same-language performance. In fact, the gap in most scenarios is more pronounced for NDCG@20 and MRR@20 compared to Hit@20.
Benchmark
User
Lang.
Doc
Lang.
Hit
@20
NDCG
@20
MRR
@20
Legal arabic arabic 92.4 65.9 60.6
arabic english 89.8 60.6 53.9
english arabic 55.8 29.9 22.9
english english 86.5 58.7 52.0
same-lang. 89.4 62.3 56.3
cross-lang. 72.8 45.3 38.4
Travel arabic arabic 92.8 85.5 84.4
arabic english 91.0 71.7 66.8
english arabic 80.0 62.5 58.1
english english 93.6 88.3 87.8
same-lang. 93.2 86.9 86.1
cross-lang. 85.5 67.1 62.5
(a) BGE-M3 embedder
Benchmark
User
Lang.
Doc
Lang.
Hit
@20
NDCG
@20
MRR
@20
Legal arabic arabic 87.5 66.4 64.0
arabic english 50.6 27.4 21.5
english arabic 40.7 21.4 16.1
english english 87.7 62.8 57.7
same-lang. 87.6 64.6 60.8
cross-lang. 45.6 24.4 18.8
Travel arabic arabic 90.0 82.8 81.5
arabic english 54.3 37.4 33.3
english arabic 63.9 35.1 26.9
english english 94.9 89.5 89.1
same-lang. 92.5 86.2 85.3
cross-lang. 59.1 36.3 30.1
(b) M-E5 embedder
Table 5: Retriever Performance across language combinations. Retriever performance is reported using three metrics: Hit@20, NDCG@20, and MRR@20. Results are presented for each embedder, benchmark and for each of the four possible user–document language combinations. In addition, we report same-language and cross-language scores, defined as the mean scores over combinations where the user and document languages match or differ, respectively.
(a) Legal benchmark – BGE-M3 embedder
(b) Legal benchmark – M-E5 embedder
(c) Travel benchmark – BGE-M3 embedder
(d) Travel benchmark – M-E5 embedder
Figure 2: Retrieval NDCG@20 scores across benchmarks and embedders. Each figure corresponds to a specific combination of benchmark and embedding. Bars represent retrieval NDCG@20 scores in percentages, with 95% confidence intervals shown as black error lines. Different retrieval policies are distinguished by color and texture. Results are grouped by benchmark segments defined by the user-document language combination, as well as the overall benchmark retrieval performance.
(a) Legal benchmark – BGE-M3 embedder
(b) Legal benchmark – M-E5 embedder
(c) Travel benchmark – BGE-M3 embedder
(d) Travel benchmark – M-E5 embedder
Figure 3: Retrieval MRR@20 scores across benchmarks and embedders. Each figure corresponds to a specific combination of benchmark and embedding. Bars represent retrieval MRR@20 scores in percentages, with 95% confidence intervals shown as black error lines. Different retrieval policies are distinguished by color and texture. Results are grouped by benchmark segments defined by the user-document language combination, as well as the overall benchmark retrieval performance.
A.5 Imbalanced corpora
In this section, we explore imbalanced corpora where one language dominates. We added experiments with different bilingual corpus ratios to examine whether the observed trends persist.
Given a target fraction $X$ of English documents, we construct a corpus in the same manner as the original one in the paper, but retain the English version of each document with probability $X$. The existing corpus corresponds to $X=50\%$, and we added two new corpora for $25\%$ and $75\%$. We evaluated three methods: (i) the direct approach, (ii) the balanced approach that retrieves 10 documents from each language, denoted Balanced-Equal, and (iii) a new method we call Balanced-Weighted, which retrieves English and Arabic documents in proportion to their ratio in the corpus (e.g., the top-$5$ documents in English and the top-$15$ in Arabic for $X=25\%$). The results are presented in Table 6.
User-Doc Langs. Retriever 25% English 50% English 75% English
Same-lang. Direct 95±1% 93±2% 92±2%
Same-lang. Balanced (Equal) 95±1% 95±2% 93±1%
Same-lang. Balanced (Weighted) 96±1% 95±2% 93±1%
Cross-lang. Direct 79±2% 86±3% 86±2%
Cross-lang. Balanced (Equal) 89±2% 93±2% 94±2%
Cross-lang. Balanced (Weighted) 90±2% 93±2% 92±2%
(a) BGE-M3 Embedder
User-Doc Langs. Retriever 25% English 50% English 75% English
Same-lang. Direct 96±1% 92±2% 92±2%
Same-lang. Balanced-Equal 96±1% 96±2% 94±1%
Same-lang. Balanced-Weighted 96±1% 96±2% 94±1%
Cross-lang. Direct 65±3% 59±4% 56±3%
Cross-lang. Balanced-Equal 91±2% 94±2% 93±2%
Cross-lang. Balanced-Weighted 93±2% 94±2% 92±2%
(b) M-E5 Embedder
Table 6: Hit@20 scores across different corpus imbalances for the Travel benchmark
We draw 2 notable conclusions from Table 6: (1) The Balanced-Equal baseline is stable in its performance across the 3 corpora, up to statistical noise. (2) The improvement of Balanced-Equal compared to the Direct baseline, as well as its competitiveness when compared to Balanced-Weighted remain across all settings, including the imbalanced corpora.